# CodeGRAG: Extracting Composed Syntax Graphs for Retrieval Augmented Cross-Lingual Code Generation

### v1

@@ Abstract
Utilizing large language models to generate codes has shown promising meaning in software development revolution. Despite the intelligence shown by the general large language models, their specificity in code generation can still be improved due to the syntactic gap and mismatched vocabulary existing among natural language and different programming languages. In addition, programming languages are inherently logical and complex, making them hard to be correctly generated. Existing methods rely on multiple prompts to the large language model to explore better solutions, which is expensive. In this paper, we propose Syntax Graph Retrieval Augmented Code Generation (CodeGRAG) to enhance the performance of LLMs in single-round code generation tasks. CodeGRAG extracts and summarizes the control flow and data flow of code blocks to fill the gap between programming languages and natural language. The extracted external structural knowledge models the inherent flows of code blocks, which can facilitate LLMs for better understanding of code syntax and serve as a bridge among different programming languages. CodeGRAG significantly improves the code generation ability of LLMs and can even offer performance gain for cross-lingual code generation, e.g., C++ for Python.
@@ 1 Introduction
Despite the success in natural language understanding and generation, there is still a gap between the natural language and programming language as illustrated in Figure 1. Code, unlike natural language, possesses a well-defined structure that governs its syntax and semantics. This structure provides valuable information about the relationships between different parts of the code, the flow of execution, and the overall organization of the functions Jiang et al. (2021); Guo et al. (2020). General-purpose LLMs regard a code block as a sequence of tokens. By ignoring the inherent structure of codes, they miss out on essential cues that could help them better understand and generate code. In addition, the multi-lingual code generation abilities of LLMs remain a challenge due to the gap between different programming languages. Moreover, the training data of different programming languages can be unbalanced, which leads to a biased ability of multi-lingual code generation.
@@
Previous methods on code representation learning seek to interpret code blocks based on the abstract syntax tree, data flow graph (Aho et al., 2006), control flow graph (Allen, 1970), and the intermediate states (Lattner and Adve, 2004), where the semantic level and logical level information of codes are better captured. These methods gain the performance improvement in code classification (Long et al., 2022), code search (Zeng et al., 2023), etc, showcasing the benefits of capturing the structures of codes. In this paper, we aim to inject the structural knowledge of codes into the general-purpose large language models for a better understanding of codes. By understanding the underlying structure of code, the LLMs can 1) extract the inherent semantic and logic flow of codes for higher quality generation and 2) transfer knowledge across different programming languages, which enables cross-lingual code generation.
However, tuning the base model for structural code understanding is expensive and heavy. Recently, retrieval augmented generation (RAG) has shown great performance gains in facilitating LLMs on complex and temporal-related tasks, well utilizing the in-context-learning ability to instruct LLMs for enhanced generation. In this paper, we propose to develop a code RAG framework to retrieve relevant structural code representations to fill the gap between natural language and programming language, provide hints for enhanced code generation, and utilize the structure information of codes to offer cross-lingual code generation augmentation.
The main challenges of designing an effective framework of RAG for code generation can be summarized in threefolds. 1) How to design effective structural views of code blocks to represent the inherent syntax of codes, which could fill the gap between natural language and programming languages, inform LLMs of external syntactic knowledge, and model the underlying cross-lingual knowledge. 2) How to measure the structural similarities between codes and how to align the natural language and programming language for joint similarity computation. 3) How to obtain an informative query format for efficient knowledge querying, which can offer consistent useful information extraction under different sets of problems that are diversified in syntax and styles.
Concretely, we propose to 1) represent the code blocks with the composed syntax graph based on the data-flow-graph and control-flow-graph, which extracts both the semantic level and the logical level information of code blocks, 2) use hybrid-GNN and pretrained cross-lingual code search models to measure the distances, and 3) obtain the problem description and the function declaration as the query content.
@@
  • We propose an efficient and effective method named CodeGRAG that bridges the gap among natural language and programming languages, transfers knowledge among different programming languages, and utilizes the in-context learning ability of LLMs for enhanced code generation. The generation process requires only one calling of LLMs and can offer cross-lingual enhancement.
  • We propose an effective graphical view to purify the semantic and logic knowledge from the code space, which offers more useful information than the raw code block and can summarize the cross-lingual knowledge.
  • We propose an effective retrieval method to retrieve relevant code knowledge for generation enhancement, where the cross-modality alignment and structure preserving are achieved.
2 Related Work
Figure 2: The illustration of the overall process of CodeGRAG.
Code Search. The code search methods can be summarized into three folds. Early methods utilizes sparse search to match the query and codes (Hill et al., 2011; Yang and Huang, 2017), which suffers from mismatched vocabulary due to the gap between natural language and codes. Neural methods (Cambronero et al., 2019; Gu et al., 2021) then focus on mapping the query and codes into a joint representation space for more accurate retrieval. With the success of pretrained language models, many methods propose to use pretraining tasks to improve the code understanding abilities and align different language spaces. For example, CodeBERT (Feng et al., 2020) is pretrained on NL-PL pairs of 6 programming languages with the masked language modeling and replaced token detection task. CodeT5 (Wang et al., 2021) supports both code-related understanding and generation tasks through bimodal dual generation. UniXcoder (Guo et al., 2022) integrates the aforementioned pretraining tasks, which is a unified cross-modal pre-trained model.
@@ 3.1 Overview
In this paper, we leverage both generative models and retrieval models to produce results that are both coherent and informed by the external syntax knowledge. The overall process of CodeGRAG is illustrated in Figure 2, which consists of three stages as follows.
Syntax Knowledge Base Preparation. We analyze the syntax and control information of code blocks and extract their graphical views to better represent the codes. This process can be formulated as, $\forall c_{i}\in D_{\text{pool}}$:
@@
where $c_{i}$ is the raw code block and $g_{i}$ is the extracted graphical view of the corresponding code block.
Knowledge Querying. Given a target problem to be completed, we generate informative query of it and use it to retrieve knowledge from the constructed knowledge base. The process can be formulated as:
@@
Knowledge Augmented Generation. After obtaining the knowledge index, we access the corresponding entry of the knowledge base and obtain the graphical syntax knowledge to inform the LLMs for enhanced code generation. The process can be formulated as:
$c_{i},g_{i}$ $\xleftarrow{}\text{KB}[i],$ (5)
result $\xleftarrow{}\text{LLM}(p,g_{i}).$ (6)
3.2 Syntax Graph Knowledge Base
In this section, we discuss how to extract informative views for code blocks and how to summarize the retrieved knowledge. An illustration is given in Figure 3. To capture both the semantic and the logical information, we propose to combine the data flow graph (Aho et al., 2006) and the control flow graph (Allen, 1970) with the read-write signals (Long et al., 2022) to represent the code blocks, both of them are constructed on the base of the abstract syntax tree.
After obtaining the composed syntax graphs for code blocks in the retrieval pool, we use them for retrieval enhancement, where both the semantic level and the logic level information are preserved.
3.3 Knowledge Querying
The main problems of the retrieval lie in: 1) how to fill the gap between natural languages and programming language and the gap among different programming languages; 2) how to properly measure the similarity of code blocks, e.g., the unique structure existing in programming languages; 3) how to design the informative query content.
3.3.1 Query Content
@@ 3.3.2 Retrieval Function
To expressively represent each of the components, we use the pretrained NL2Code model encoder $\phi_{1}(\cdot)$ to represent the problem description and code snippets and design a graph neural network $\phi_{2}(\cdot)$ to represent the graphical view. The retrieval function can be formulated as:
$\mathbf{h^{V}}$ $=\psi(\phi_{1}(V_{c}),\phi_{2}(V_{g})),$ (7)
$\mathbf{h^{Q}}$ $=\phi_{1}(Q_{p}\|Q_{c}),$ (8)
@@
We use models pretrained to align natural language and programming language to serve as the encoder $\phi_{1}$:
  • CodeT5+ (Wang et al., 2023) is an advanced encoder-decoder foundation LLM, tailored for diverse code understanding and generation tasks. It is designed for various code-related tasks, capable of operating in different modes (encoder-only, decoder-only, encoder-decoder). It employs a range of pretraining tasks, including span denoising, causal language modeling on code, and text-code contrastive learning, matching and CLM tasks. This approach improves its performance in retrieval tasks by effectively aligning text and code data.
  • UnixCoder (Guo et al., 2022) is a cross-modal pre-trained model optimized for both code understanding and generation tasks. Leveraging a multi-layer Transformer structure with mask attention matrices, it outperforms traditional unified encoder-decoder models in auto-regressive tasks like code completion. UniXcoder uniquely employs multi-modal inputs, including code comments and abstract syntax trees (AST), enhancing its code representation capabilities. It employs three types of language modeling tasks: masked language modeling, unidirectional language modeling, and denoising objectives. Additionally, it introduces two advanced pre-training tasks: multi-modal contrastive learning and cross-modal generation, which refine the semantic representation of code fragments.
And we design a graph neural network $\phi_{2}$ to preserve the semantic and logical information of code blocks. The representation of each node $\mathbf{n}_{i}^{(0)}$ and edge $\mathbf{e}_{i}^{(0)}$ are first initialized with vectors corresponding to the node text and edge text encoded by $\phi_{1}$. A message passing process is first conducted to fuse the semantic and structural information into each node representation.
@@
The different encoding networks are optimized via the contrastive learning based self-supervised training, which includes the intra-modality contrastive learning and inter-modality contrastive learning. The intra-modality constrastive learning serves for preserving the modality information, while the inter-modality contrastive learning serves for modality alignment.
Alignment Contrastive Learning. There are two types of alignment to be ensured during the retrieval process: 1) Question-Answer (QA) alignment and 2) Code-Graph (CG) alignment. We define the positive pairs for QA alignment purpose as $\mathcal{I}^{+}_{QA}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{i}^{Q}\rangle|i\in D_{\text{train}}\}$ and define the negative pairs for QA alignment purpose as $\mathcal{I}^{-}_{QA}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{j}^{Q}\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$
@@
3.4 Knowledge Augmented Generation
After we obtain the summarized retrieved knowledge, we append it to the prompt text for knowledge augmented generation. The prompt used is illustrated in Figure 4.
Figure 4: The illustration of the used prompt.
The prompt consists of three components: the system prompt illustrated in the blue part, the retrieved knowledge and hints illustrated in the green part, and the problem (including task description, function declaration, etc.) illustrated in the yellow part. The three parts are concatenated to be fed into LLMs for knowledge augmented generation.
Table 1: Main results of Pass@1 on Humaneval-X.
Model Size Model Humaneval-X C++ Humaneval-X Python
6B GPT-J 7.54 11.10
6B CodeGen-Multi 11.44 19.41
6.7B InCoder 9.50 16.41
13B CodeGeeX 17.06 22.89
16B CodeGen-Multi 18.05 19.22
16B CodeGen-Mono 19.51 29.28
15B StarCoder 31.55 32.93
15B WizardCoder 29.27 57.3
15B Pangu-Coder2 45.12 64.63
- GPT-3.5-Turbo 57.93 71.95
- GPT-3.5-Turbo + Code-RAG 60.37 72.56
- GPT-3.5-Turbo + CodeGRAG 62.20 72.56
- GPT-3.5-Turbo + Cross-Lingual Code-RAG 62.20 70.12
- GPT-3.5-Turbo + Cross-Lingual CodeGRAG 64.02 77.44
@@ 4 Experiments
In this section, we conduct experiments to answer the following research questions (RQs).
@@
    Does the proposed CodeGRAG model the high-level thought-of-codes? Can CodeGRAG offer cross-lingual augmentation?
@@
    How does the retrieval algorithms impact the performance?
@@ 4.1 Setup
In this paper, we evaluate CodeGRAG with the widely used HumanEval-X (Zheng et al., 2023) dataset, which is a multi-lingual code benchmark. We use greedy decoding strategy to do the generation. The evaluation metric is Pass@1.
We evaluate the multi-lingual code generation abilities of 1) models with less than 10 billion parameters: GPT-J (Radford et al., 2023), CodeGen-Multi (Nijkamp et al., 2022), and InCoder(Fried et al., 2022); 2) models with 10-20 billion parameters: CodeGeeX (Zheng et al., 2023), CodeGen-Multi (Nijkamp et al., 2022), CodeGen-Mono (Nijkamp et al., 2022), StarCoder (Li et al., 2023), WizardCoder (Luo et al., 2023), and Pangu-Coder2 (Shen et al., 2023); 3) close-sourced GPT-3.5 model. The detailed statistics of the model sizes are displayed in Table 1.
@@ 4.2 Main Results
The main results are summarized in Table 1. Code-RAG represents directly using the raw code block as the retrieved knowledge and CodeGRAG refers to using the graphical view as the knowledge. From the results, we can draw the following conclusions:
RQ1. The proposed CodeGRAG could offer performance gain against the base model, which validates the effectiveness of the proposed RAG for code generation framework.
RQ2. The model with CodeGRAG could beat model with raw code block augmented generation. From the results, we can see that the proposed graph view could summarize the useful structural syntax information and filter out the noises, which could offer more informative knowledge hints than the raw code block.
RQ3. Python code generation is improved with cross-lingual retrieval, which uses C++ code and corresponding graphical views as the retrieved knowledge for an enhanced generation. Therefore, the proposed CodeGRAG could offer cross-lingual augmentation. This validates that the proposed graphical view could abstract the high-level and inherent information (e.g., the control and data flow to solve a specific problem) of the code blocks, which are shared across different programming languages regarding solving the same problem.
Table 2: Results with different content encoders (yielding different retrieval models).
Content Encoder $\phi_{1}(\cdot)$ Method Humaneval-X C++ Humaneval-X Python
NA GPT-3.5-Turbo 57.93 71.95
CodeT5 + Code-RAG 60.37 72.56
+ CodeGRAG 62.20 72.56
+ Cross-Lingual Code-RAG 61.59 70.12
+ Cross-Lingual CodeGRAG 62.20 76.83
UnixCoder + Code-RAG 58.54 71.95
+ CodeGRAG 62.20 71.95
+ Cross-Lingual Code-RAG 62.20 66.46
+ Cross-Lingual CodeGRAG 64.02 77.44
4.3 Impact of Retrieval Methods (RQ4)
In this section, we explore how different retrieval methods influence the final performance. We use CodeT5 and UnixCoder as the text and code encoder, which are both trained with multi-lingual data and dedicated for multi-lingual search. The results are summarized in Table 2.
From the results, we could see that the augmentation performance can differ with different retrieval content encoders. Specifically, the cross-lingual augmentation performance increases with UnixCoder being the encoder compared with that of CodeT5 being encoder. Since UnixCoder involves in the abstract syntax tree for structure capturing, while CodeT5 does not. This validates the effectiveness of involving structure information. In addition, we can see that with better cross-lingual retrieval accuracy, directly using raw codes as cross-lingual hints will bias the model generation, while feeding graphical views of syntax will boost the model performance.

### v2

# CodeGRAG: Bridging the Gap between Natural Language and Programming Language via Graphical Retrieval Augmented Generation

@@ Abstract
Utilizing large language models to generate codes has shown promising meaning in software development revolution. Despite the intelligence shown by the general large language models, their specificity in code generation can still be improved due to the syntactic gap and mismatched vocabulary existing among natural language and different programming languages. In this paper, we propose CodeGRAG, a Graphical Retrieval Augmented Code Generation framework to enhance the performance of LLMs. CodeGRAG builds the graphical view of code blocks based on the control flow and data flow of them to fill the gap between programming languages and natural language, which can facilitate natural language based LLMs for better understanding of code syntax and serve as a bridge among different programming languages. To take the extracted structural knowledge into the foundation models, we propose 1) a hard meta-graph prompt template to transform the challenging graphical representation into informative knowledge for tuning-free models and 2) a soft prompting technique that injects the domain knowledge of programming languages into the model parameters via finetuning the models with the help of a pretrained GNN expert model. Various experiments and ablations are done on four datasets including both the C++ and python languages to validate the hard meta-graph prompt, the soft prompting technique, and the effectiveness of the objectives for pretrained GNN expert. CodeGRAG improves the code generation ability of LLMs and can even offer performance gain for cross-lingual code generation. The implementation is available at https://anonymous.4open.science/r/Code-5970/ .
@@ 1 Introduction
However, syntactic gap and mismatched vocabulary among natural language (NL) and programming languages (PL) exist, hindering LLM’s performance on code generation. As illustrated in Figure 1, programming language (marked in blue) contains special tokens such as “int” or “++” that natural language (marked in yellow) doesn’t possess, leading to vocabulary mismatch. Besides, the relations between tokens in programming languages are often structural, e.g., the complex branching and jumps, whereas natural language is arranged simply in sequential manner, leading to syntactic gap. For example, in the control flow graph of the raw code (marked in pink), two “if” blocks (marked in purple) are adjacent and are executed sequentially under certain condition, but they appear to be intervaled in raw textual code.
As discussed above, the innate structures of programming languages are different from that of the sequential-based natural language. The challenges of enhancing a general-purposed large language models for code-related tasks can be summarized into two folds.
@@
(C1) How to solve the gap between different languages and better interpret the inherent logic of code blocks. Code, unlike natural language, possesses a well-defined structure that governs its syntax and semantics. This structure provides valuable information about the relationships between different parts of the code, the flow of execution, and the overall organization of the functions Jiang et al. (2021); Guo et al. (2020). General-purpose LLMs regard a code block as a sequence of tokens. By ignoring the inherent structure of codes, they miss out on essential cues that could help them better understand and generate code. In addition, the multi-lingual code generation abilities of LLMs is challenging due to the gap among different programming languages.
(C2) How to inject the innate knowledge of programming languages into general purpose large language models for enhancement. Despite the well representation of the programming knowledge, the ways to inject the knowledge into the NL-based foundation models is also challenging. The structural representation of codes could be hard to understand, which poses a challenge to the capability of the foundation models.
To solve the above challenges, we propose CodeGRAG, a graphical retrieval augmented generation framework for code generation. For (C1), we propose to interpret the code blocks using the composed graph based on the data-flow and control-flow of the code block, which extracts both the semantic level and the logical level information of the code. The composed graphical view could 1) better capture the innate structural knowledge of codes for NL-based language models to understand and 2) model the innate function of code blocks that bridging different programming languages. For (C2), we propose a meta-graph prompting technique for tuning-free models and a soft-prompting technique for tuned models. The meta-graph prompt summarizes the overall information of the extracted graphical view and transforms the challenging and noisy graphical representation into informative knowledge. The soft-prompting technique deals with the graphical view of codes with a pretrained GNN expert network and inject the processed knowledge embedding into the parameters of the general-purpose foundation models with the help of supervised finetuning.
@@
  • Novel GraphRAG framework for code generation. We propose CodeGRAG that bridges the gap among natural language and programming languages, transfers knowledge among different programming languages, and enhances the ability of LLMs for code generation. CodeGRAG requires only one calling of LLMs and can offer multi-lingual enhancement.
  • Effective graphical view to inform and stimulate the structural programming knowledge of LLMs. We propose an effective graphical view to purify the semantic and logic knowledge from the code space, which offers more useful information than the raw code block and can summarize the cross-lingual knowledge.
  • Effective soft prompting technique to preserve the programming domain knowledge and inject it into LLMs parameters. We propose an effective soft prompting technique, which injects the domain knowledge of programming languages into the model parameters via finetuning LLMs with the assistance of a pretrained GNN expert model.
@@ 2 Methodology
Figure 2: Overview of CodeGRAG. (Top) Knowledge Preparation. We extract the control flow and data flow of each external code block and compose them using the read-write signal to obtain the semantic and logical expression of each code block, which is then abstracted into graphical view as hard knowledge document and embedded into GraphEmb as soft knowledge document. The GraphEmb is encoded by a pretrained GNN expert model constrained by the alignment and structure preserving objectives. (Bottem) Retrieval Augmented Generation. We extract query from the task input and retrieve from the external corpus. For tuning free models, we use the hard graphical view to stimulate the structural programming knowledge of LLMs for enhanced generation. For tunable models, we use the soft GraphEmb and inject the programming domain knowledge into LLMs parameters via finetuning them with the GNN expert signals. The expert signals informed LLMs can then produce enhanced generation.
@@ 2.1 Overview
In this paper, we leverage both generative models and retrieval models to produce results that are both coherent and informed by the expert graphical knowledge of programming language. The overall process of CodeGRAG is illustrated in Figure 2, which mainly consists of three stages: graphical knowledge base preparation, knowledge querying, and graphical knowledge augmented generation.
2.2 Graphical Knowledge Base Preparation
In this section, we discuss how to extract informative graphical views for code blocks. We analyze the syntax and control information of code blocks and extract their graphical views to better represent the codes. This process can be formulated as, $\forall c_{i}\in D_{\text{pool}}$:
@@ 2.2 Graphical Knowledge Base Preparation
where $c_{i}$ is the raw code block and $g_{i}$ is the corresponding extracted graphical view.
To capture both the semantic and the logical information, we propose to combine the data flow graph (Aho et al., 2006) and the control flow graph (Allen, 1970) with the read-write signals (Long et al., 2022) to represent the code blocks, both of them are constructed on the base of the abstract syntax tree.
An illustration of the composed graphical view is given in Figure 3. After obtaining the composed syntax graphs for code blocks in the retrieval pool, we use them to inform the general-purpose LLMs to bridge the gap between NL and PLs, where both the semantic level and the logic level information are preserved.
2.3 Knowledge Querying
Given a target problem to be completed, we generate informative query of it and use it to retrieve graphical knowledge from the constructed knowledge base. The process can be formulated as:
@@ 2.3 Knowledge Querying
The main problems of the retrieval lie in: 1) how to design the informative query content and 2) how to align the different modalities.
2.3.1 Query Extractor
@@ 2.3.2 Retriever
To expressively represent the components, we use the encoder $\phi(\cdot)$ of the pretrained NL2Code model to represent the problem description and code snippets. The retrieval function is:
$\mathbf{h^{V}}$ $=\phi(V_{c}\|V_{g}),$ (5)
$\mathbf{h^{Q}}$ $=\phi(Q_{p}\|Q_{c}),$ (6)
@@
2.4 Graphical Knowledge Augmented Generation
After we obtain the returned graphical view, we inject it to the foundation LLMs for graphical knowledge augmented generation. Since the graphical view is hard to understand, we propose 1) a meta-graph template to transform the graphical view into informative knowledge for tuning-free model and 2) a soft prompting technique to tune the foundation models for their better understanding of the graphical views with the assistance of an expert GNN model.
2.4.1 Hard Meta-Graph Prompt
The original graphical view of a code block could contain hundreds of nodes and edges. A full description of it could cost a overly long context, along with the understanding challenge posed by the long edge lists. Therefore, we propose to use a meta-graph template to abstract the information of the graphical view. The abstracted meta-graph consists of the canonical edge types and node types, which describes the basic topology of the graphical view (Sun & Han, 2013), with the textual features obtained from the ASTs contained in the node and edge features.
Then we use the meta-graph template to transform the retrieved graphical view into digestable knowledge and insert it into the final prompt for generation. As illustrated in Figure 5, the final prompt consists of three components: the system prompt illustrated in the blue part, the retrieved knowledge and hints illustrated in the green part, and the problem (including task description, function declaration, etc.) illustrated in the yellow part. The three parts are concatenated to be fed into LLMs for knowledge augmented generation.
Figure 4: Hard meta-graph prompt.
Figure 5: Soft prompting.
2.4.2 Soft Prompting with the Expert
Directly hard prompt to the LLMs poses a challenge to the digesting capability of the backbone LLMs, which could fail under the case where the backbone LLMs cannot well understand the graph components.
To compress the graphical knowledge into model parameters and help the backbone LLMs to better understand the programming language, we propose a soft prompting technique. The overall procedure can summarized into expert encoding of graphical views, finetuning with the expert signal, and inference.
Expert Encoding of Graphical Views. We design a graph neural network to preserve the semantic and logical information of code blocks. The representation of each node $\mathbf{n}_{i}^{(0)}$ and edge $\mathbf{e}_{i}^{(0)}$ are first initialized with vectors corresponding to the node text and edge text encoded by $\phi_{1}$. A message passing process is first conducted to fuse the semantic and structural information into each node representation.
@@ 2.4.2 Soft Prompting with the Expert
The expert encoding network is optimized via the contrastive learning based self-supervised training, which includes the intra-modality contrastive learning and inter-modality contrastive learning. The intra-modality constrastive learning serves for preserving the modality information, while the inter-modality contrastive learning serves for modality alignment.
  • Alignment Contrastive Learning. There are two types of alignment to be ensured: 1) NL-Code (NC) alignment and 2) Code-Graph (CG) alignment. We define the positive pairs for NC alignment purpose as $\mathcal{I}^{+}_{NC}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{i}^{Q}\rangle|i\in D_{\text{train}}\}$ and define the negative pairs for NC alignment purpose as $\mathcal{I}^{-}_{NC}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{j}^{Q}\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$
@@
Finetuning with the Expert Soft Signal. To help the backbone LLMs to digest the graphical views, we tune the LLMs with the expert soft signal using supervised finetuning. The prompt for finetuning consists of the system prompt, retrieved knowledge where the expert encoded graphical view is contained using a token embedding, and task prompt, which is illustrated in Figure 5.
Inference. After the finetuning stage, we used the tuned models to generate codes using the soft prompting template as illustrated in Figure 5.
@@ 3 Experiments
    Can soft prompting enhance the capability of the backbone LLMs? Does finetuning with the soft prompting outperforms the simple supervised finetuning?
@@
    Are the proposed pretraining objectives for the GNN expert effective?
@@ 3.1 Setup
In this paper, we evaluate CodeGRAG with the widely used HumanEval-X (Zheng et al., 2023) dataset, which is a multi-lingual code benchmark and CodeForce dataset in which we collect real-world programming problems from codeforces [^1] https://codeforces.com/ website. For CodeForce dataset we include problems categorized by different difficulty levels corresponding to the website and select $469$ problems of difficulty level A for testing. We use greedy decoding strategy to do the generation. The evaluation metric is Pass@1.
For the size of retrieval pool, we use 11,913 C++ code snippets and 2,359 python code snippets. Due to the limited access, we do not use a large retrieval corpus for our experiment, which can be enlarged by other people for better performance. We also attach the graph extraction codes for both languages and all other expeirment codes here: https://anonymous.4open.science/r/Code-5970/
For the fintuning details, the learning rate and weight decay for the expert GNN training is 0.001 and 1e-5, repectively. We apply 8-bit quantization and use LoRA for parameter-efficient fine-tuning. The rank of the low-rank matrices in LoRA is uniformly set to 8, alpha set to 16, and dropout is set to 0.05. The LoRA modules are uniformly applied to the Q and V parameter matrices of the attention modules in each layer of the LLM. All the three models are optimized using the AdamW optimizer. For the CodeContest dataset, totally 10609 datapoints are used, and for APPS dataset, 8691 data samples are used to train the model.
@@ 3.2 Main Results
The main results are summarized in Table 1 and Table 2.
Table 1: Results of Hard Meta-Graph Prompt on Humaneval-X. (Pass@1)
Model Retrieved Knowledge C++ Python
GPT-3.5-Turbo N/A 57.93 71.95
Code Block (Nashid et al., 2023; Lu et al., 2022) 60.37 72.56
Meta-Graph 62.20 72.56
(Multi-Lingual) Code-Block (Nashid et al., 2023; Lu et al., 2022) 62.20 70.12
(Multi-Lingual) Meta-Graph 64.02 77.44
GPT-4omini N/A 65.24 78.66
Code Block (Nashid et al., 2023; Lu et al., 2022) 65.24 78.66
Meta-Graph 65.85 79.88
(Multi-Lingual) Code-Block (Nashid et al., 2023; Lu et al., 2022) 65.85 79.27
(Multi-Lingual) Meta-Graph 67.07 80.49
Table 2: Results of Soft Prompting. (Pass@1)
Model Finetune CodeForce (C++) APPS (Python)
Gemma 7b N/A 12.83 5.09
SFT 14.76 21.09
Soft Prompting 19.13 26.15
Llama2 13b N/A 9.61 7.29
SFT 11.88 12.06
Soft Prompting 13.62 12.74
CodeLlama 7b N/A 5.20 24.41
SFT 9.87 26.15
Soft Prompting 11.09 30.26
From the results, we can draw the following conclusions:
RQ1. The proposed CodeGRAG could offer performance gain against the base model, which validates the effectiveness of the proposed graphical retrieval augmented generation for code generation framework.
RQ2. The model informed by the meta-graph (CodeGRAG) could beat model informed by the raw code block. From the results, we can see that the proposed graph view could summarize the useful structural syntax information and filter out the noises, which could offer more informative knowledge hints than the raw code block. In addition, inserting the intermediate representations of codes into the prompt can stimulate the corresponding programming knowledge of LLMs.
RQ3. From Table 2, we can see that finetuning with the expert soft prompting could offer more performance gain than that brought by simple supervised finetuning. This validates the effectiveness of the designed pretraining expert network and the technique of finetuning with soft prompting, which injects the programming domain knowledge into the LLMs parameters and inform the models with the structural information for gap filling.
3.3 Impacts of the pretraining objectives for the expert GNN (RQ4)
To study the effectiveness of the proposed pretraining objectives for the expert GNN, we remove each objective to yield different expert GNNs. The results are in Table 3.
Table 3: Ablation studies on the GNN pretraining losses.
Model Finetune CodeForce (C++) APPS (Python)
Gemma 7b Soft Prompting 19.13 26.15
w/o Alignment 7.88 28.58
w/o Structure-Preserving 11.70 21.50
Llama2 13b Soft Prompting 13.62 12.74
w/o Alignment 11.79 10.76
w/o Structure-Preserving 5.50 11.09
CodeLlama 7b Soft Prompting 11.09 30.26
w/o Alignment 10.92 29.45
w/o Structure-Preserving 10.66 26.59
From the results, we could see that both the Alignment and the Structure Preserving contribute to the expressiveness of the expert GNN model. The alignment pretraining objective helps to promote the alignment among natural language, programming language, and their graphical views. The structure preserving objective helps to preserve the innate data-flows and control-flows information of code blocks. The two objectives collaborate with each other to yield expressive programming domain knowledge GNN expert model, which encodes external programming knowledge and injects the knowledge into LLMs parameters.